OpenStudy (marissalovescats):

Eliminate the parameter x = 3 cos t, y = 3 sin t Hi an explanation on how to go about this would be much appreciated!

OpenStudy (anonymous):

recall the famous identity \(\cos^2(t)+\sin^2(t)=1\)

OpenStudy (marissalovescats):

2 of my favorite OpenStudyers typing, this makes me happy! Thanks for the wanting to help @satellite73 and @vishweshshrimali5 And yes I know there's something I need to do with that identity but I'm unsure

OpenStudy (anonymous):

do it!

OpenStudy (anonymous):

no dear, just square and add do it write \[x^2+y^2=...\] and you will see the \(t\) magically disappear

OpenStudy (anonymous):

i made a typo there square \(3\cos(t)\) and square \(3\sin(t)\) and add them up

OpenStudy (anonymous):

let me know what you get then we will see why you knew the answer to begin with

OpenStudy (marissalovescats):

Uhm I'm not sure because I'm not fully understanding. But if I squared both wouldn't I get 9cos^2+9sin^2t

OpenStudy (anonymous):

yes !

OpenStudy (marissalovescats):

Question: Why are we even adding them?

OpenStudy (anonymous):

that means \[x^2+y^2=9\cos^2(x)+9\sin^2(x)\] factor the \(9\) from the right hand side

OpenStudy (anonymous):

i will answer after you do it and see the \(t\) go bye bye

OpenStudy (marissalovescats):

x^2+y^2=9(1)

OpenStudy (anonymous):

yay the t is gone and you see that you have a circle of radius 3

OpenStudy (marissalovescats):

Is that eliminating the parameter? So that's all I had to do?

OpenStudy (anonymous):

you knew this before you started the problem the unit circle has points \((x,y)\) that sastifies the equation \(x^2+y^2=1\) they are also given by the points \((\cos(t), \sin(t))\) i.e. parametrized by the "angle"

OpenStudy (anonymous):

yes, that is all you have to do you had the parameter \(t\) in \(x=3\cos(t), y = 3\sin(t)\) now you have \[x^2+y^2=9\] the parameter is gone

OpenStudy (anonymous):

if it as \[x=\cos(t),y=\sin(t)\] it is the unit circle multiplying by 3 makes the radius 3 instead of 1

OpenStudy (marissalovescats):

Oh okay, thank you :)

OpenStudy (vishweshshrimali5):

I was going to reply the same thing but @satellite73 stole my answer :P Thanks for thinking of so highly of me @marissalovescats

OpenStudy (anonymous):

yw with a little practice you will do this in your head i realize you might think at the beginning that there is some one method for "eliminating the parameter" but there is not really one way to do it

OpenStudy (marissalovescats):

I think highly of all!

OpenStudy (vishweshshrimali5):

That's a good thing but never underestimate your self. So, got the answer ?

OpenStudy (marissalovescats):

Yup, I did!

OpenStudy (vishweshshrimali5):

 This is the graphical solution :)